Solve by substitution
4x+3y=-5
x=y-3

Answer by funmath(2933)

Solve by substitution
E1) 4x+3y=-5
E2) x=y-3
Subtitute y-3 in for x in E1 and solve for y:
4(y-3)+3y=-5
4y-12+3y=-5
7y-12=-5
7y-12+12=-5+12
7y=7
7y/7=7/7
y=1
Substitute 1 in for y in E2 and solve for x:
x=1-3
x=-2
The solution (x,y)=(-2,1)
Happy Calculating!!!